OpenStudy (anonymous):

Find all solutions to the equation interval [0, 2pi) 2cosx sinx - cos x=0

OpenStudy (anonymous):

divide the equation by cos x, u wil get : 2sin x - 1 = 0 so, sin x= 1/2 thus, x= arcsin (1/2). u can solve it.